# VLP: Vision Language Planning for Autonomous Driving

### v3

@@ Abstract
Autonomous driving is a complex and challenging task that aims at safe motion planning through scene understanding and reasoning. While vision-only autonomous driving methods have recently achieved notable performance, through enhanced scene understanding, several key issues, including lack of reasoning, low generalization performance and long-tail scenarios, still need to be addressed. In this paper, we present VLP, a novel Vision-Language-Planning framework that exploits language models to bridge the gap between linguistic understanding and autonomous driving. VLP enhances autonomous driving systems by strengthening both the source memory foundation and the self-driving car’s contextual understanding. VLP achieves state-of-the-art end-to-end planning performance on the challenging NuScenes dataset by achieving 35.9% and 60.5% reduction in terms of average L2 error and collision rates, respectively, compared to the previous best method. Moreover, VLP shows improved performance in challenging long-tail scenarios and strong generalization capabilities when faced with new urban environments. Code will be made public upon publications.
@@ 1 Introduction
The way humans process and interpret visual driving scenes naturally involve a coherent cognitive framework that maintains a consistent logical flow. This enables humans effortlessly make correct decisions even when faced with previously unseen data or scenarios, eliminating generalization issues in driving tasks. Then, this raises a pertinent question: How can ADS achieve human-like driving, leveraging the same principles that enable humans to navigate diverse and dynamic road environments?
@@
Recent advancements in large-language models (LLMs), which are trained on vast and diverse textual datasets, have led to unprecedented common-sense ability and generalization performance across unseen data and tasks for the natural language processing domain [5, 25, 8]. The superior common-sense capability of LLMs have led to the emergence of multi-modal LLMs for diverse applications ranging from medical imaging to robotics [33, 9, 4]. However, incorporation of the reasoning ability of LLMs into real-world autonomous driving tasks, to address generalization and long-tail scenarios, is yet to be fully-explored.
@@
To bridge this gap, we propose a Vision Language Planning (VLP) framework, which integrates the common-sense capability of LLMs into vision-based ADS for safe self-driving. Our VLP consists of two key components: Agent-centric Learning Paradigm (ALP) and Self-driving-car-centric Learning Paradigm (SLP), leveraging LLMs to enhance the ADS from reasoning and decision-making aspects, respectively.
@@
The BEV feature map serves as the source memory pool in ADS for downstream decoding tasks. It summarizes and encodes the driving environment surrounding the self-driving car, including vehicles, pedestrians, lanes, and more, into a unified feature map. Hence, capturing comprehensive and necessary details in each local position of BEV is critical for safe and precise self-driving performance. To enhance the local semantic representation and reasoning capabilities of BEV, we introduce an innovative Agent-centric Learning Paradigm (ALP) module. ALP integrates the consistent feature space of a pretrained language model to revamp the agent features on the BEV, actively shaping semantics, and guiding the BEV reasoning process. Leveraging on the common sense and logic flow embedded in the language model, our ALP equips the ADS with robustness and consistent BEV feature space, enhancing its effectiveness in diverse driving scenarios.
@@
In ADS, the planning module aggregates information from the preceding perception and prediction phases to make the final decisions for self-driving. This global perspective culminates in the formation of a planning query, directly influencing the safety and accuracy of the self-driving navigation. Considering the critical role of the planning module within ADS, we also present a novel Self-driving-car-centric Learning Paradigm (SLP) to elevate the decoding and acquiring information ability of the planning query. In the SLP, we align the planning query with intended goals and the ego-vehicle driving status by leveraging the knowledge encoded in the pretrained language model. The language model’s comprehension capabilities contribute to more informed decision-making during the planning phase as well as enabling a more robust planning query formation process.
@@
Through VLP, we bridge the gap between human-like reasoning and autonomous driving, enhancing the model’s contextual awareness and its ability to generalize effectively in complex, ever-changing real-world scenarios. The main contributions of this work are summarized as follows:
@@
  • We propose VLP, a Vision Language Planning model, which incorporates reasoning capability of LLMs into vision-based autonomous driving systems as an enhancement of motion planning and self-driving safety.
@@
  • To the best of our knowledge, this is the first work introducing LLMs into multiple stages of ADS to address the generalization ability in new cities and long-tail cases.
@@ 2.3 Autonomous Driving with Language Models
In recent years, numerous works have been proposed for extending LLMs to the embodied AI domain to improve zero-shot generalization performance [35, 4, 9, 30]. These works incorporate sensory inputs from embodied agents in the form of language with vision and language data for decision making process [9, 4]. While these methods have mainly focused on the robotics domain, few works have been proposed for leveraging embodied LLMs for autonomous driving tasks [32, 24, 11, 34, 31]. In particular, DiLu [32] and GPT-Driver [24] propose GPT-based driver agents for closed-loop simulation tasks. In [31], an open loop driving commentator, which combines vision and low-level driving actions with language, is proposed to interpret driving actions and for reasoning. Our work significantly differs from these approaches by leveraging LLMs for end-to-end motion planning task for real-world driving applications.
@@ 3 Methodology
Figure 2: a) The overview of proposed vision language planning (VLP) framework. VLP enhances ADS from self-driving BEV-reasoning and self-driving decision-making aspects, through two innovative modules, ALP and SLP, respectively. Leveraging LLM and contrastive learning, ALP conducts agent-wise learning for refining local details on BEV, while SLP engages sample-wise learning for advancing global context understanding ability of the ADS. VLP is only active during training, ensuring no additional parameters or computations are introduced during inference. b) Prompt formats used in VLP.
We present a Vision Language Planning (VLP) model, which equips the ADS with the capacity to emulate human common sense and engage in contextual reasoning for safe motion planning. Our proposed VLP model, illustrated in Fig. 2a, comprises two innovative components that leverage LLMs in both local and global contexts. The Agent-centric Learning Paradigm (ALP) concentrates on refining local details to enhance source memory reasoning, while the Self-driving-car-centric Learning Paradigm (SLP) focuses on guiding the planning process for the self-driving-car (SDC). Below, we first provide an overview of how general ADS work in Sec.3.1. Subsequently, we describe our proposed ALP and SLP in detail in Sec. 3.2 and Sec. 3.3, respectively.
@@ 3.2 Agent-centric Learning Paradigm: Enhancing BEV Source Memory
The BEV feature map serves as the foundational source memory pool and holds a pivotal role in ADS. Ensuring that a BEV map provides comprehensive information with intricate details for driving is essential for enabling downstream decoders to make safe, precise, and human-like decisions. However, the BEV in ADS is derived from multi-view camera images rather than a true bird’s-eye-view image, which can introduce discrepancies between the produced BEV and the expected BEV representation. Therefore, to mitigate the discrepancies, we introduce the Agent-centric Learning Paradigm (ALP) to align the produced BEV with a true bird’s-eye-view map. Through direct supervision on the BEV map with LLM, our ALP enables the refinement of local details and alignment with the desired human perspective.
@@
Agent expectation features. To ensure that local BEV features express the desired information, we conduct a BEV-expectation alignment process by leveraging LLM and contrastive learning. We precisely define the perceptual information expected from the corresponding agent, such as agent labels, bounding boxes, and future trajectories. These driving-related ground-truth information, which should also be embedded in the local BEV feature is formulated into a prompt as illustrated in Fig. 2b. The description is then passed to the language encoder, $LM$, to generate the corresponding agent expectation feature. We apply an MLP layer $\mathcal{F}_{bev}$ to adapt the expectation feature to the BEV feature space. Then, the agent expectation features are concatenated along the batch to generate an Agent-Expectation tensor denoted as $\mathcal{A}_{exp}\in\mathbb{R}^{N_{B}\times C}$. The procedure can be formulated as in Eq. (2):
@@ 4 Experiments
We present the implementation details of our experiments, including the dataset and training details, in Sec. 4.1, and validate the effectiveness of our approach for open-loop planning and perception / prediction in Sec. 4.3. Then, we evaluate the generalization ability of our model in Sec.4.4. Ablation studies are provided in Sec. 4.5. Qualitative results are presented in the supplementary material.
@@ 4.1 Implementation Details
Training. We incorporate our proposed VLP training strategy into vision-only approaches of UniAD [16] and VAD [18], and refer to this enriched versions as VLP-UniAD and VLP-VAD, respectively, for clarity when reporting the results. All methods are trained with the same hyper-parameters reported in the respective baselines [16, 18] for commensurate comparison. For language model, we have used language model in CLIP [27]. In VLP study, we set T=6 (3 seconds) for planning, which is same as baselines [16] [18]. Experiments are conducted with 8 NVIDIA Tesla A100 GPUs. More details regarding experimental setup can be found in supplementary materials.

### v4

@@ Abstract
Autonomous driving is a complex and challenging task that aims at safe motion planning through scene understanding and reasoning. While vision-only autonomous driving methods have recently achieved notable performance, through enhanced scene understanding, several key issues, including lack of reasoning, low generalization performance and long-tail scenarios, still need to be addressed. In this paper, we present VLP, a novel Vision-Language-Planning framework that exploits language models to bridge the gap between linguistic understanding and autonomous driving. VLP enhances autonomous driving systems by strengthening both the source memory foundation and the self-driving car’s contextual understanding. VLP achieves state-of-the-art end-to-end planning performance on the challenging NuScenes dataset by achieving 35.9% and 60.5% reduction in terms of average L2 error and collision rates, respectively, compared to the previous best method. Moreover, VLP shows improved performance in challenging long-tail scenarios and strong generalization capabilities when faced with new urban environments.
@@ 1 Introduction
The way humans process and interpret visual driving scenes naturally involve a coherent cognitive framework that maintains a consistent logical flow. This enables humans effortlessly make correct decisions even when faced with previously unseen data or scenarios, eliminating generalization issues in driving tasks. Then, this raises a pertinent question: How can ADS achieve human-like driving to navigate diverse and dynamic road environments?
@@
Recent advancements in language models (LMs) have led to unprecedented common-sense ability and generalization performance across unseen data and tasks for the natural language processing domain [5, 25, 8]. The superior common-sense capability of LMs have led to the emergence of multi-modal LMs for diverse applications ranging from medical imaging to robotics [34, 9, 4]. However, incorporation of the reasoning ability of LMs into real-world autonomous driving tasks, to address generalization and long-tail scenarios, is yet to be fully-explored.
@@
To bridge this gap, we propose a Vision Language Planning (VLP) framework, which integrates the common-sense capability of LMs into vision-based ADS for safe self-driving. Our VLP consists of two key components: Agent-centric Learning Paradigm (ALP) and Self-driving-car-centric Learning Paradigm (SLP), leveraging LMs to enhance the ADS from reasoning and decision-making aspects, respectively.
@@
The BEV feature map serves as the source memory pool in ADS for downstream decoding tasks. It summarizes and encodes the driving environment surrounding the self-driving car, including vehicles, pedestrians, lanes, and more, into a unified feature map. Hence, capturing comprehensive and necessary details in each local position of BEV is critical for safe and precise self-driving performance. To enhance the local semantic representation and reasoning capabilities of BEV, we introduce an innovative Agent-centric Learning Paradigm (ALP) module. ALP integrates the consistent feature space of a pretrained LM to revamp the agent features on the BEV, actively shaping semantics, and guiding the BEV reasoning process. Leveraging on the common sense and logic flow embedded in the LM, our ALP equips the ADS with robustness and consistent BEV feature space, enhancing its effectiveness in diverse driving scenarios.
@@
In ADS, the planning module aggregates information from the preceding perception and prediction phases to make the final decisions for self-driving. This global perspective culminates in the formation of a planning query, directly influencing the safety and accuracy of the self-driving navigation. Considering the critical role of the planning module within ADS, we also present a novel Self-driving-car-centric Learning Paradigm (SLP) to elevate the decoding and acquiring information ability of the planning query. In the SLP, we align the planning query with intended goals and the ego-vehicle driving status by leveraging the knowledge encoded in the pretrained LM. The language model’s comprehension capabilities contribute to more informed decision-making during the planning phase as well as enabling a more robust planning query formation process.
@@
Through VLP, we bridge the gap between human-like reasoning and autonomous driving, enhancing the model’s contextual awareness and its ability to generalize effectively in complex, ever-changing real-world scenarios, as shown in Fig.1. The main contributions of this work are summarized as follows:
@@
  • We propose VLP, a Vision Language Planning model, which incorporates reasoning capability of LMs into vision-based autonomous driving systems as an enhancement of motion planning and self-driving safety.
@@
  • To the best of our knowledge, this is the first work introducing LM into multiple stages of ADS to address the generalization ability in new cities and long-tail cases.
@@ 2.3 Autonomous Driving with Language Models
In recent years, numerous works have been proposed for extending LMS to the embodied AI domain to improve zero-shot generalization performance [36, 4, 9, 31]. These works incorporate sensory inputs from embodied agents in the form of language with vision and language data for decision making process [9, 4]. While these methods have mainly focused on the robotics domain, few works have been proposed for leveraging embodied LMs for autonomous driving tasks [33, 24, 11, 35, 32, 26]. In particular, DiLu [33] and GPT-Driver [24] propose GPT-based driver agents for closed-loop simulation tasks. In [32], an open loop driving commentator, which combines vision and low-level driving actions with language, is proposed to interpret driving actions and for reasoning. Our work significantly differs from these approaches by leveraging LMs for end-to-end motion planning task for real-world driving applications.
@@ 3 Methodology
Figure 2: a) The overview of proposed vision language planning (VLP) framework. VLP enhances ADS from self-driving BEV-reasoning and self-driving decision-making aspects, through two innovative modules, ALP and SLP, respectively. Leveraging LM and contrastive learning, ALP conducts agent-wise learning for refining local details on BEV, while SLP engages sample-wise learning for advancing global context understanding ability of the ADS. VLP is only active during training, ensuring no additional parameters or computations are introduced during inference. b) Prompt formats used in VLP.
We present a Vision Language Planning (VLP) model, which equips the ADS with the capacity to emulate human common sense and engage in contextual reasoning for safe motion planning. Our proposed VLP model, illustrated in Fig. 2a, comprises two innovative components that leverage LMs in both local and global contexts. The Agent-centric Learning Paradigm (ALP) concentrates on refining local details to enhance source memory reasoning, while the Self-driving-car-centric Learning Paradigm (SLP) focuses on guiding the planning process for the self-driving-car (SDC). Below, we first provide an overview of how general ADS work in Sec.3.1. Subsequently, we describe our proposed ALP and SLP in detail in Sec. 3.2 and Sec. 3.3, respectively.
@@ 3.2 Agent-centric Learning Paradigm: Enhancing BEV Source Memory
The BEV feature map serves as the foundational source memory pool and holds a pivotal role in ADS. Ensuring that a BEV map provides comprehensive information with intricate details for driving is essential for enabling downstream decoders to make safe, precise, and human-like decisions. However, the BEV in ADS is derived from multi-view camera images rather than a true bird’s-eye-view image, which can introduce discrepancies between the produced BEV and the expected BEV representation. Therefore, to mitigate the discrepancies, we introduce the Agent-centric Learning Paradigm (ALP) to align the produced BEV with a true bird’s-eye-view map. Through direct supervision on the BEV map with LM, our ALP enables the refinement of local details and alignment with the desired human perspective.
@@
Agent expectation features. To ensure that local BEV features express the desired information, we conduct a BEV-expectation alignment process by leveraging LM and contrastive learning. We precisely define the perceptual information expected from the corresponding agent, such as agent labels, bounding boxes, and future trajectories. These driving-related ground-truth information, which should also be embedded in the local BEV feature is formulated into a prompt as illustrated in Fig. 2b. The description is then passed to the language encoder, $LM$, to generate the corresponding agent expectation feature. We apply an MLP layer $\mathcal{F}_{bev}$ to adapt the expectation feature to the BEV feature space. Then, the agent expectation features are concatenated along the batch to generate an Agent-Expectation tensor denoted as $\mathcal{A}_{exp}\in\mathbb{R}^{N_{B}\times C}$. The procedure can be formulated as in Eq. (2):
@@ 4.1 Implementation Details
Training. We incorporate our proposed VLP training strategy into vision-only approaches of UniAD [16] and VAD [18], and refer to this enriched versions as VLP-UniAD and VLP-VAD, respectively, for clarity when reporting the results. All methods are trained with the same hyper-parameters reported in the respective baselines [16, 18] for commensurate comparison. For language model, we use the pretrained LM in CLIP [28]. In VLP study, we set T=6 (3 seconds) for planning, which is same as baselines [16] [18]. Experiments are conducted with 8 NVIDIA Tesla A100 GPUs. More details regarding experimental setup can be found in supplementary materials.
@@ 5 Conclusion
[…]
Supplementary Material
6 Experiment Setup
Model Components. UniAD is composed of a BEV extractor followed by five transformer decoder-based P3 modules for 3D object tracking, mapping, motion forecasting, occupancy prediction, and planning, respectively. VAD is composed of a BEV extractor followed by four transformer decoder-based modules for 3D object detection, scene segmentation, motion forecasting, and planning, respectively.
Model Hyperparameters. We provide the detailed training configurations for UniAD-based and VAD-based models in the Tab. 11.
Configs UniAD-based VAD-based
point cloud range $[-51.2,-51.2,-5.0,51.2,51.2,3.0]$ $[-15.0,-30.0,-2.0,15.0,30.0,2.0]$
transformer decoder dimension 256 256
BEV size $200\times 200$ $100\times 100$
queue length 3 3
motion predict steps 12 6
motion predict modes 6 6
planning steps 6 6
backbone RN101 RN50
optimizer AdamW AdamW
learning rate 2e-4 2e-4
weight decay 0.01 0.01
epoch 20 60
batch size 8 8
Table 11: Training configurations for UniAD-based and VAD-based models.
7 Ablation Study
Prompt Information. We present ablation studies on prompt information for UniAD and VAD, accompanied by downstream task results in Tab. 15 and 16. Notably, training with completed ground truth information embedded in the prompts consistently yields superior performance across all downstream tasks for both UniAD and VAD. In comparison to the baseline (1st line in each table), models equipped with VLP, regardless of the ground truth information included, consistently outperform the model without VLP. These observations underscore the efficacy of the proposed VLP in enhancing model performance.
Different LMs. We explored several pretrained LMs for integration with our VLP. As shown in Tab. 12, each integration led to consistent improvements in decision-making over two baselines, proving the effectiveness of VLP design (leverages LM locally and globally). Also, it underscores our contribution of integrating language understanding into visual models for autonomous driving, irrespective of the specific choice of the LM. Our choice was driven by a balance between computational efficiency and effectiveness. Integration with LLAMA will be our future work.
ALP and SLP. As shown in Tab. 13, both SLP and ALP consistently improve all tasks. SLP focuses more on the final planning as it works for the ego-car query while ALP works for the shared BEV map. When ALP and SLP operate concurrently, the model puts more efforts on optimizing the final planning. While there is a relatively smaller margin of improvement in scene segmentation and occupancy prediction, the VLP still significantly surpasses the baseline (+VLP: +5.9%). Trajectory planning is the most important step which guarantees the safety and efficiency of ADS. Improvement on it is the main target of VLP.
Pretrained LM Planning Pretrained LM Planning
avg.L2 $\downarrow$ avg.Col $\downarrow$ avg.L2 $\downarrow$ avg.Col $\downarrow$
UniAD VAD
- 1.03 0.31 - 0.82 0.93
+VLP +VLP
GPT2 0.75 (+27.2%) 0.16 (+48.4%) GPT2 0.61 (+25.6%) 0.29 (+68.8%)
CLIP-RN50x64-LM 0.74 (+28.2%) 0.16 (+48.4%) CLIP-RN50x64-LM 0.62 (+24.4%) 0.30 (+67.7%)
CLIP-RN101-LM 0.76 (+26.2%) 0.21 (+32.3%) CLIP-RN101-LM 0.57 (+30.5%) 0.35 (+62.4%)
CLIP-ViT-L/14-336px-LM 0.73 (+29.1%) 0.24 (+22.6%) CLIP-ViT-L/14-336px-LM 0.64 (+22.0%) 0.45 (+51.6%)
Table 12: Open-loop planning results with various LMs.
VLP Vectorized Scene Seg. Occupancy Pred. Plan Val. Plan Singapore
[…]
- 39.8 63.4  40.2  54.7  33.5 0.78 0.38 0.78 0.39
+SLP 47.2 64.2  40.7  55.8  34.5 0.50 0.23 0.66 0.25
+ALP 47.6 64.5  41.0  56.1  34.7 0.52 0.26 0.68 0.28
+VLP 45.7 64.1  40.2  55.9  34.1 0.55 0.15 0.63 0.20
Table 13: Ablation study on each component of VLP.
[…]
VAD [18] 17.6 29.9 0.79 0.33 0.95 0.65 0.17
VLP-VAD $\checkmark$ 19.1 31.2 0.76 0.33 1.00 0.50 0.25
VLP-VAD $\checkmark$ $\checkmark$ 20.4 33.2 0.75 0.33 0.83 0.60 0.19
Table 14: 3D Object detection in challenging long-tail scenarios.
Included Ground Truth Tracking Mapping Motion Forecasting Occupancy Prediction Planning
Label Bbox Trajectory Command AMOTA↑ AMOTP↓ IDS↓ IoU-lane↑ IoU-road↑ minADE↓ minFDE↓ MR↓ IoU-n.↑ IoU-f.↑ VPQ-n.↑ VPQ-f.↑ avg.L2↓ avg.Col.↓
35.2 1.353 720 29.5 66.7 0.83 1.24 0.187 59.5 38.5 49.9 29.4 0.93 0.38
$\checkmark$ $\checkmark$ $\checkmark$ 35.5 1.331 670 30.8 66.8 0.78 1.23 0.180 59.8 39.1 51.2 30.7 0.87 0.31
$\checkmark$ $\checkmark$ $\checkmark$ 35.9 1.331 650 31.2 67.5 0.75 1.19 0.162 61.9 38.5 53.1 31.4 0.82 0.32
$\checkmark$ $\checkmark$ $\checkmark$ 35.4 1.340 710 31.1 67.2 0.81 1.20 0.179 62.1 38.7 53.8 31.5 0.89 0.36
$\checkmark$ $\checkmark$ $\checkmark$ 36.1 1.329 610 31.6 68.9 0.76 1.12 0.168 61.7 38.9 54.1 31.9 0.80 0.29
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 36.2 1.320 620 31.7 69.1 0.72 1.11 0.156 62.4 39.3 54.2 32.9 0.78 0.24
Table 15: Detailed ablations on the effectiveness of prompt information with UniAD-based models. Our ablation studies on prompt information for UniAD reveal a significant performance boost when training with completed ground truth information embedded in the prompts.
Included Ground Truth 3D Object Detection Vectorized Scene Segmentation Motion Forecasting Planning
Label Bbox Trajectory Command mAP↑ NDS↑ Boundary↑ Divider↑ Crossing↑ mIoU↑ minADE↓ minFDE↓ MR↓ EPA↑ avg.L2↓ avg.Col.↓
22.6 33.3 43.3 43.6 34.0 40.3 0.86 1.20 0.143 0.526 0.98 0.80
$\checkmark$ $\checkmark$ $\checkmark$ 24.9 35.7 45.2 42.5 32.9 40.9 0.76 1.06 0.127 0.522 0.64 0.23
$\checkmark$ $\checkmark$ $\checkmark$ 25.7 37.6 46.5 44.4 30.4 40.4 0.78 1.05 0.119 0.549 0.56 0.30
$\checkmark$ $\checkmark$ $\checkmark$ 27.0 38.7 44.9 40.0 34.8 40.0 0.86 1.19 0.141 0.532 0.59 0.36
$\checkmark$ $\checkmark$ $\checkmark$ 26.6 38.0 44.0 40.3 31.4 38.9 0.79 1.06 0.121 0.551 0.59 0.26
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 27.3 39.0 47.1 45.3 36.0 42.8 0.77 1.05 0.117 0.551 0.52 0.17
Table 16: Detailed ablations on the effectiveness of prompt information with VAD-based models. Our ablation studies on prompt information for VAD shows a considerable performance improvement when training with completed ground truth information embedded in the prompts.
Pretrained LM Tracking Mapping Motion Forecasting Occupancy Prediction Planning
AMOTA↑ AMOTP↓ IDS↓ IoU-lane↑ IoU-road↑ minADE↓ minFDE↓ MR↓ IoU-n.↑ IoU-f.↑ VPQ-n.↑ VPQ-f.↑ avg.L2↓ avg.Col.↓
- 35.2 1.353 720 29.5 66.7 0.83 1.24 0.187 59.5 38.5 49.9 29.4 0.93 0.38
CLIP-RN50x64-LM 32.9 1.384 780 29.2 65.4 0.70 1.01 0.140 59.7 38.0 50.3 29.1 0.81 0.23
CLIP-RN101-LM 36.2 1.320 620 31.7 69.1 0.72 1.11 0.156 62.4 39.3 54.2 32.9 0.78 0.24
CLIP-ViT-L/14-336px-LM 32.6 1.358 923 29.8 67.1 0.70 1.01 0.138 59.0 37.6 49.1 28.4 0.57 0.87
GPT-2 34.1 1.379 784 29.8 67.0 0.72 1.03 0.148 60.6 38.3 50.3 28.9 0.93 0.22
Table 17: Detailed ablations investigating the impact of various language models on UniAD-based models.
8 Long-tail Generalization for 3D Object Detection
Tab. 14 presents the generalization ability of each VLP component on long-tail cases for 3D Object Detection. The results highlight the efficacy of each component in mitigating the long-tail detection problem. Particularly, the inclusion of SLP leads to a noticeable improvement over the baseline, and the combined utilization of both SLP and ALP further enhances the generalization ability. The long-tail classes, including construction vehicles, buses, motorcycles, bicycles, and trailers, constitute approximately 6% of the nuScenes dataset.
9 Visualization
We present several qualitative comparisons with the baseline in Fig. 3-9, using green arrows to highlight areas where our model outperforms the baseline. The visual comparison illustrates that our Vision Language Planning (VLP) framework help navigate the self-driving car in a more efficient and safer way.
10 Why human-like?
Humans interpret the visual scene as contextual cognitive semantics instead of plain digits, enabling them to navigate in unseen environments and recognize rare objects. We imbue the visual system with a more human-like and intuitively expected feature space via inspiring the reasoning and decision-making processes with designed prompts and pretrained LMs. During inference, our system retains the same enriched feature space and robust capabilities, which is evidenced by the marked improvements in the new-city generalization and long-tail tracking/detection.
Figure 3: Qualitative comparison between UniAD and Ours. Green arrow is used to highlight areas where our VLP outperforms the baseline. The results indicate that our VLP enables the self-driving car to navigate more efficiently and safely.
Figure 4: Qualitative comparison between UniAD and Ours. Green arrow highlights areas where our VLP outperforms the baseline.
Figure 5: Qualitative comparison between UniAD and Ours. Green arrow highlights areas where our VLP outperforms the baseline.
Figure 6: Qualitative comparison between UniAD and Ours. Green arrow highlights areas where our VLP outperforms the baseline.
Figure 7: Qualitative comparison between UniAD and Ours. Green arrow highlights areas where our VLP outperforms the baseline.
Figure 8: Qualitative comparison between UniAD and Ours. Green arrow highlights areas where our VLP outperforms the baseline.
Figure 9: Qualitative comparison between UniAD and Ours. Green arrow highlights areas where our VLP outperforms the baseline.